Username12:

A punch recipe calls for 2/3 of a pint of fruit juice for each pint of soda. The ratio of soda to fruit juice in the punch is __. to __.

KIllerwolf:

6:4 ig

Katrin369:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @KIllerwolf 6:4 ig \(\color{#0cbb34}{\text{End of Quote}}\) Next time don't give the direct answer.

KIllerwolf:

@Katrin369 srry

Username12:

Thank you so much!!!

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @KIllerwolf 6:4 ig \(\color{#0cbb34}{\text{End of Quote}}\) 6/4 or simplified by 2 result 3/2

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @KIllerwolf 6:4 ig \(\color{#0cbb34}{\text{End of Quote}}\) 6/4 or simplified by 2 result 3/2 \(\color{#0cbb34}{\text{End of Quote}}\) what is \[\left( \frac{ 2 }{ 3 } \right)^{-1} = \frac{ 1 }{ \frac{ 2 }{ 3 } }= \frac{ 3 }{ 2 }\]

jhonyy9:

@AZ

Username12:

Thank you :)